Suppose that the Price level = 140, the Supply of Money = $20 billion, and the Velocity of Money is 35. What does real GDP equal?

Select one:

a. $4 billion

b. $5 billion

c. $8 billion

d. $10 billion

e. $12 billion
Answer #1

Velocity of money=GDP/money supply

So GDP =velocity of money×money supply

So GDP=35×20=700

Real GDP =nominal GDP/deflator

Real GDP=700/140

=$5 billion.

So option B is the correct statement.